# E2E-MFD: Towards End-to-End Synchronous Multimodal Fusion Detection

### v2

@@ 4.1 Dataset and Implementation Details
We conduct experiments on four widely-used visible-infrared image datasets: TNO [50], RoadScene [25], M3FD [16] and DroneVehicle [2]. TNO and RoadScene are just used to evaluate MF performance. M3FD is adopted to evaluate both MF and OD performance. Besides, DroneVehicle is utilized for oriented OD. RoadScene with 37 image pairs, TNO with 42 image pairs and M3FD with 300 pairs are only used for the MF task in the testing stage, and the MF network is trained by the M3FD dataset which is divided into a training set (2,940 image pairs) and a testing set (1,260 image pairs). We conduct all the experiments with one GeForce RTX 3090 GPU, and the code of M3FD is based on Detectron2 [51], while the code of DroneVehicle is based on MMDetection 2.26.0 [52] and MMRotate 0.3.4 [53]. The pretrained DiffusionDet is used for the initialization of the OD network. In the training phase, E2E-MFD is optimized by AdamW with a batch size of 1. We set the learning rate to $2.5\times 10^{-5}$ and the weight decay as $41e-$ 4. The default training iteration is only 15,000. Different setting information about DroneVehicle can be found in the Appendix A.4 and the implementation of metric can be found in the Appendix A.1.
@@ 4.2 Main Results
Results on Multimodal Image Fusion. Qualitative results of different fusion methods are depicted in Figure 3. All the fusion methods can fuse the main features of the infrared and visible images to some extent and we can observe two remarkable advantages of our method. First, the significant characteristics of infrared images can be effectively highlighted by our method. As shown in Figure 3, our M3FD fusion image captures the person riding a motorcycle. In comparison with other methods, our method demonstrates high contrast and recognition of the objects. Second, our method preserves rich details from the visible images, including color and texture. In Figure 3, our advantages are evident in the fusion images across the M3FD dataset, such as the clear outline of the white car’s rear and a man on a motorcycle. While retaining a substantial amount of detail, our method maintains a high resolution without introducing blurriness. In contrast, other methods fail to achieve these two advantages simultaneously. Sequentially, we provide quantitative results of different fusion methods in Table 1. Our E2E-MFD generally achieves the best metric values. Specifically, the largest average value on MI proves that our method transfers more considerable information from both source images. Values of EN reveal that our results contain edged details and the highest contrast between objects and the background. The large VIF shows our fusion results have high-quality visual effects and small distortion compared with the source images. Moreover, our method achieves the fastest training time to finish the joint learning at one stage which means that faster iterative updates can be done on new datasets. The test time to generate a fused image ranked third. For more visualization results, see Appendix A.5.
@@
Results on Multimodal Object Detection. To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. We follow the same parameter settings, and the visualization results are shown in Figure 4. The detection results are poor when using only single-modal image inputs, with instances of missed detection, such as the motorcycle and rider next to the car and people on the far right in the image. Almost all fusion methods reduce missed detection and improve confidence by fusing information from both modalities. Through the design of an end-to-end fusion detection synchronous optimization strategy, we obtain fusion images that are visually and detection-friendly, especially for occluded and overlapping objects, as seen in the blue ellipse with the motorcycle and the overlapping people on the far right in the image. To further assess the quality of fusion images, we conduct a fair comparison between our method and the SOTA methods on YOLOv5s. As shown in Table 2, the SOTA methods demonstrate a performance improvement compared to single-modal detection, indicating that well-fused images can effectively assist downstream tasks. In contrast, the fusion image we generate has achieved the best performance on YOLOv5s. Additionally, the performance of fusion images on DiffusionDet is also impressive, albeit slightly lower than when optimizing fusion and detection tasks simultaneously. Thanks to the collaborative optimization of both tasks, the detection performance is further enhanced. Furthermore, even when compared to end-to-end object detection methods, our approach demonstrates significant performance improvements. This better underscores the advantages of our training paradigm and the effectiveness of our method. For further visualization results, please refer to the Appendix A.6.
Results on Multimodal Oriented Object Detection. As shown in Table 3, our fusion detection synchronous optimization strategy achieves the highest accuracy. Furthermore, the outstanding detection performance on YOLOv5s-OBB [56] and LSKNet [62] using the generated fusion images (with at least 5.7% and 3.1% higher AP values compared to single modalities) demonstrate the robustness of our method. This validates the superior quality of the fusion images, indicating that they are not only visually appealing but also provide rich information for the detection task. For further visualization results, please refer to the Appendix A.4.
@@ 4.3 Ablation Studies
Effect of Gradient Matrix Task-Alignment. To verify the effectiveness of GMTA, we compare separate optimizations for MF and OD, as well as joint optimization, w/o and w indicating whether to use GMTA. Specifically, MF represents using only $\mathcal{L}_{u}$ to optimize the fusion network, OD represents using only $\mathcal{L}_{d}$ to optimize the object detection network, and E2E-MFD represents simultaneous optimization of the fusion and detection networks using both $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ loss functions. The results, as shown in Table 4, indicate that methods incorporating GMTA optimization constraints with shared weights achieve the best results for both MF and OD. This is because MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. Therefore, optimizing the entire network with shared loss functions may be influenced by local optimal solutions of individual tasks.The accuracy of E2E-MFD (w/o GMTA) shows a slight decrease compared to separately training the detection network. In contrast, GMTA orthogonalizes the gradients of shared parameters corresponding to the two tasks, allowing the joint network to converge to an optimal point with fusion detection weights. For more comparison of different multi-task Learning, please refer to Appendix A.3.
@@
Study of Branches in the Object-Region-Pixel Phylogenetic Tree. We investigate the combination of the pixel feature mining module (0) and the region feature refinement module (1,2,3,4). The results are shown in Table 5. It can be observed that with the increase in the number of branches, the fusion network achieves higher image fusion quality and object detection performance. Region features provide the fusion network with multi-level semantic features of objects. However, when higher-level semantic object information is added, the performance of the fusion network declines. This is because the detailed information contained in the deeper layers of the backbone structure shared with the detection network decreases abruptly, which may affect the fusion network’s absorption of these features, thereby influencing pixel-level fusion tasks. The different feature map visualization can be seen in Appendix A.2.
@@ 5 Conclusion
Within this paper, synchronous-parallel task-alignment optimization is proposed to formulate fusion and detection in a harmonious one-stage training process. We introduce a phylogenetic tree structure to simulate these two tasks in different visual perceptions needed for device task requirements. In addition, we align the orthogonal components of the fusion detection linear system of the gradients. By unrolling the model to a well-designed fusion network and diffusion-based detection network, we can generate a visual-friendly result for fusion and object detection in an efficient way without tedious training steps and inherent optimization barriers.
@@ A.1 Metric
A.2 Visialization of Feature Maps
Figure 6: Feature map visualization of various branches in the ORPPT.
A.3 More Ablation Studies
To compare the effectiveness of different Multi-task learning (MTL) methods on our algorithm, we selected three robust MTL optimization methods. As shown in Table 6, all MTL methods addressed the conflict between the MF and OD tasks to varying degrees. By introducing the concept of Gradient Matrix Task-Alignment (GMTA), we achieved a better balance in the gradient optimization process between the two tasks, resulting in the best performance across all metrics.
Table 6: Ablation of different MTL methods on M3FD.
Method EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
E2E-MFD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
PCGradYu et al. [2020] 6.13 15.01 1.48 90.59 62.71
CAGradLiu et al. [2021c] 6.17 15.05 1.48 90.71 62.74
Nash-MTLNavon et al. [2022] 6.29 15.28 1.51 90.91 62.97
E2E-MFD (w GMTA) 6.36 15.47 1.65 91.80 63.83
A.4 Experiments on DroneVehicle
DroneVehicle comprises aerial RGB-IR images captured by drones, encompassing various scenes from an aerial perspective. It contains five categories of target objects. The dataset consists of 28,439 pairs of images, divided into training, validation, and test sets. We train the MF network on the training set and perform inference directly on the test set. Finally, fusion images are generated for both the training and test sets. All detection accuracies are obtained on the test set. The pretrained LSKNet is used for the initialization of the object detection network, and we fine-tune it for 12 epochs with a batch size of 4. To fully demonstrate the effectiveness of our method, we compare some state-of-the-art (SOTA) models, and experiments are also conducted on YOLOv5s-OBB using the generated fusion images, all images are resized to 640$\times$640 and trained from scratch for 50 epochs with a batch size of 64.
The visualization results of the fusion and detection of DroneVehicle are shown in Figure 7 and 8. In the process of generating fusion images, we first obtain RGB images, then convert them to HSV format, and only take the V channel as the final fusion image. It preserves the brightness information of the objects while containing rich details from visible images. As shown in Figure 7, the first row and the second row are images extracted from the training set and the test set, respectively. Although the style of the fusion images resembles that of the infrared images, by combining visible image information, the fusion images can differentiate different parts of each object, reflecting varying degrees of brightness information. This is advantageous for fine-grained classification. Additionally, as shown in Figure 8, the visually friendly fusion images also assist in the object detection task, allowing for sufficient detection even in cases of small and dense objects.

### v4

@@ 4.1 Dataset and Implementation Details
We conduct experiments on four widely-used visible-infrared image datasets: TNO [53], RoadScene [26], M3FD [17] and DroneVehicle [3]. TNO and RoadScene are just used to evaluate MF performance. M3FD is adopted to evaluate both MF and OD performance. RoadScene with 37 image pairs, TNO with 42 image pairs and M3FD with 300 pairs are only used for the MF task in the testing stage, and the MF network is trained by the M3FD dataset which is divided into a training set (2,940 image pairs) and a testing set (1,260 image pairs). Besides, DroneVehicle consists of 28,439 image pairs is utilized to train and test MF and OD for oriented objects. We conduct all the experiments with one GeForce RTX 3090 GPU, and the code of M3FD is based on Detectron2 [54], while the code of DroneVehicle is based on MMDetection 2.26.0 [55] and MMRotate 0.3.4 [56]. On the M3FD dataset, the pretrained DiffusionDet is used for the initialization of the OD network. In the training phase, E2E-MFD is optimized by AdamW with a batch size of 1. We set the learning rate to $52.5e-$ 5 and the weight decay as $41e-$ 4. The default training iteration is only 15,000. On the DroneVehicle dataset, the pretrained LSKNet [57] is used for the initialization of the object detection network, and we fine-tune it for 12 epochs with a batch size of 4. The E2E-MFD is optimized by AdamW and the learning rate and the weight decay is set to $41e-$ 4 and 0.05.
@@ 4.2 Main Results
Results on Multimodal Image Fusion. Qualitative results of different fusion methods are depicted in Figure 3. All the fusion methods can fuse the main features of the infrared and visible images to some extent and we can observe two remarkable advantages of our method. First, the significant characteristics of infrared images can be effectively highlighted by our method. Our M3FD fusion image captures the person riding a motorcycle. In comparison with other methods, our method demonstrates high contrast and recognition of the objects. Second, our method preserves rich details from the visible images, including color and texture. Our advantages are evident in the fusion images across the M3FD dataset, such as the clear outline of the white car’s rear and a man on a motorcycle. While retaining a substantial amount of detail, our method maintains a high resolution without introducing blurriness. In contrast, other methods fail to achieve these two advantages simultaneously. Sequentially, we provide quantitative results of different fusion methods in Table 1. Our E2E-MFD generally achieves the best metric values. Specifically, the largest average value on MI proves that our method transfers more considerable information from both source images. Values of EN reveal that our results contain edged details and the highest contrast between objects and the background. The large VIF shows our fusion results have high-quality visual effects and small distortion compared with the source images. Moreover, our method achieves the fastest training time to finish the joint learning at one stage which means that faster iterative updates can be done on new datasets. The test time to generate a fused image ranked third.
@@
Results on Multimodal Object Detection. To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. We follow the same parameter settings, and the visualization results are shown in Figure 4. The detection results are poor when using only single-modal image inputs, with instances of missed detection, such as the motorcycle and rider next to the car and people on the far right in the image. Almost all fusion methods reduce missed detection and improve confidence by fusing information from both modalities. Through the design of an end-to-end fusion detection synchronous optimization strategy, we obtain fusion images that are visually and detection-friendly, especially for occluded and overlapping objects, as seen in the blue ellipse with the motorcycle and the overlapping people on the far right in the image. To further assess the quality of fusion images, we conduct a fair comparison between our method and the SOTA methods on YOLOv5s. As shown in Table 2, the MF methods demonstrate a performance improvement compared to single-modal detection, indicating that well-fused images can effectively assist downstream tasks. In contrast, the fusion image we generate has achieved the best performance on YOLOv5s. Additionally, the detection performance of fusion images on DiffusionDet is also impressive, albeit slightly lower than when optimizing fusion and detection tasks simultaneously with E2E-MFD. Thanks to the collaborative optimization of both tasks, the detection performance is further enhanced. Furthermore, even when compared to end-to-end object detection methods (E2E-OD), our approach demonstrates significant performance improvements. This better underscores the advantages of our training paradigm and the effectiveness of our method.
Results on Multimodal Oriented Object Detection. As shown in Table 3, our fusion detection synchronous optimization strategy achieves the highest accuracy. Furthermore, the outstanding detection performance on YOLOv5s-OBB [60] and LSKNet using the generated fusion images (with at least 5.7% and 3.1% higher AP values compared to single modalities) demonstrate the robustness of our method. This validates the superior quality of the fusion images, indicating that they are not only visually appealing but also provide rich information for the detection task.
@@ 4.3 Ablation Studies
Effect of Gradient Matrix Task-Alignment. To verify the effectiveness of GMTA, we compare separate optimizations for MF and OD, as well as joint optimization, w/o and w indicating whether to use GMTA. Specifically, MF represents using only $\mathcal{L}_{u}$ to optimize the fusion network, OD represents using only $\mathcal{L}_{d}$ to optimize the object detection network, and E2E-MFD represents simultaneous optimization of the fusion and detection networks using both $\mathcal{L}_{u}$ and $\mathcal{L}_{d}$ loss functions. The results, as shown in Table 4, indicate that methods incorporating GMTA optimization constraints with shared weights achieve the best results for both MF and OD. This is because MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. Therefore, optimizing the entire network with shared loss functions may be influenced by local optimal solutions of individual tasks. The accuracy of E2E-MFD (w/o GMTA) shows a slight decrease compared to separately training the detection network. In contrast, GMTA orthogonalizes the gradients of shared parameters corresponding to the two tasks, allowing the joint network to converge to an optimal point with fusion detection weights.
Method EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
E2E-MFD (w/o GMTA) 6.12 14.70 1.39 90.05 62.60
PCGrad[50] 6.13 15.01 1.48 90.59 62.71
CAGrad[47] 6.17 15.05 1.48 90.71 62.74
Nash-MTL[49] 6.29 15.28 1.51 90.91 62.97
E2E-MFD (w GMTA) 6.36 15.47 1.65 91.80 63.83
Table 5: Ablation of different MTL methods on M3FD.
To compare the effectiveness of different Multi-task learning methods on our algorithm, we selected three robust MTL optimization methods. As shown in Table 5, all MTL methods addressed the conflict between the MF and OD tasks to varying degrees. By introducing the concept of GMTA, we achieved a better balance in the gradient optimization process between the two tasks, resulting in the best performance.
@@
$n$ EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$
500 5.93 14.78 1.58 90.93 62.73
1000 6.36 15.47 1.65 91.80 63.83
1500 6.24 15.08 1.62 91.10 62.96
2000 6.13 14.69 1.45 90.35 62.75
Table 6: Ablation studies of the iteration parameter $n$ on M3FD dataset.
The GMTA process operates during the computation and updating stages of two gradients. The GMTA is performed approximately every $n$ iteration (gradient update), focusing on balancing the independence and coherence of various tasks. Table 6 presents the ablation analysis of the $n$ parameter. Decreasing $n$ initially disrupts task optimization due to frequent alignment, while increasing $n$ becomes crucial when the network determines task optimization directions. However, excessively large $n$ leads to significant deviations in task paths, making alignment more challenging and negatively impacting performance.
@@
Study of Branches in the Object-Region-Pixel Phylogenetic Tree. We investigate the combination of the pixel feature mining module (0) and the region feature refinement module (1,2,3,4). The results are shown in Table 7. It can be observed that with the increase in the number of branches, the fusion network achieves higher image fusion quality and object detection performance. Region features provide the fusion network with multi-level semantic features of objects. However, when higher-level semantic object information is added, the performance of the fusion network declines. This is because the detailed information contained in the deeper layers of the backbone structure shared with the detection network decreases abruptly, which may affect the fusion network’s absorption of these features, thereby influencing pixel-level fusion tasks.
Figure 6: Feature map visualization of various branches in the ORPPT.
@@ 5 Conclusion
Within this paper, an end-to-end optimization is proposed to formulate fusion and detection in a harmonious one-stage training process. We introduce a object-region-pixel phylogenetic tree structure and coarse-to-fine diffusion process to simulate these two tasks in different visual perceptions needed for diverse task requirements. In addition, we align the orthogonal components of the fusion detection linear system of the gradients by gradient matrix task-alignment. By unrolling the model to a well-designed fusion network and diffusion-based detection network, we can generate a visual-friendly result for fusion and object detection in an efficient way without tedious training steps and inherent optimization barriers.
@@ A.1 Metric
A.2 Experiment setting on YOLOv5s
To more effectively evaluate the fusion images and observe their impact on downstream detection tasks, we conduct tests using the baseline detector YOLOv5s on all SOTA methods on the M3FD dataset. All images are resized to 1024$\times$1024 and trained from scratch for 300 epochs with a batch size of 64. In addition, the experiments for DroneVehicle dataset are conducted on YOLOv5s-OBB using the generated fusion images, all images are resized to 640$\times$640 and trained from scratch for 50 epochs with a batch size of 64.
A.3 Experiments on DroneVehicle
DroneVehicle comprises aerial RGB-IR images captured by drones, encompassing various scenes from an aerial perspective. It contains five categories of target objects. The dataset consists of 28,439 pairs of images, divided into training, validation, and test sets. We train the MF network on the training set and perform inference directly on the test set. Finally, fusion images are generated for both the training and test sets. All detection accuracies are obtained on the test set. The visualization results of the fusion and detection of DroneVehicle are shown in Figure 7 and 8. In the process of generating fusion images, we first obtain RGB images, then convert them to HSV format, and only take the V channel as the final fusion image. It preserves the brightness information of the objects while containing rich details from visible images. As shown in Figure 7, the first row and the second row are images extracted from the training set and the test set, respectively. Although the style of the fusion images resembles that of the infrared images, by combining visible image information, the fusion images can differentiate different parts of each object, reflecting varying degrees of brightness information. This is advantageous for fine-grained classification. Additionally, as shown in Figure 8, the visually friendly fusion images also assist in the object detection task, allowing for sufficient detection even in cases of small and dense objects.
@@ A.3 Experiments on DroneVehicle
A.4 Experiments on Coase-to-Fine Diffusion Process
We conducted ablation experiments on CFDP in Table 8, investigating its inclusion and the number of proposed boxes. In the setting without CFDP, we maintained the backbone network while substituting CFDP with RPN (Region Proposal Network), standard components in two-stage object detectors. Results indicate that CFDP enhances detailed information capture and precise box guidance, thereby enhancing fusion image quality and detection performance. For optimal balance between performance and efficiency, we selected 500 proposal boxes.
Table 8: Ablation studies of the CFDP on M3FD dataset.
Settings Proposal boxes EN MI VIF $\text{mAP}_{50}$ $\text{mAP}_{50:95}$ Tr.Time
w/o CFDP 500 5.71 14.39 1.45 90.13 61.98 2h52m11s
w CFDP 300 6.01 14.57 1.53 90.89 63.29 2h23m45s
500 6.36 15.47 1.65 91.80 63.83 2h50m32s
1000 6.37 15.34 1.63 92.05 63.75 3h32m30s